OpenStudy (anonymous):

\[f(x) = erf\] is a function which satisfies f(1) = erf (equal approx) 0.84270 and \[f'(x) = 2 [(e ^{-x ^{2}})/(\sqrt{\pi})]\] use a linear approximation to approximate f(0.9):

OpenStudy (anonymous):

I know the equation for a linear approximenation is l(x) = f(a) + f'(a)(x-a) but I'm confused on how to insert the stuff

OpenStudy (mr.math):

It's nothing but direct substitution with x=0.9, a=1.

OpenStudy (anonymous):

Ah, I guess the only thing I'm really confused on is the f, and for f(a) do I just put 0.9 in f'(x) and whatever is the result is then f(a)?

OpenStudy (mr.math):

\(a=1 \implies f(a)=f(1)= 0.84270\).

OpenStudy (mr.math):

\[f'(a)=f'(1)=2\frac{e^{-1}}{\sqrt{\pi}}.\]

OpenStudy (mr.math):

Makes sense?

OpenStudy (anonymous):

Yeah, I get it now, thanks for your help. I appreciate it :D

OpenStudy (mr.math):

You're welcome :)